OpenStudy (anonymous):

A group of friends are addressing wedding invitations. Casey can address 21 envelopes in one hour, Anne can address 47 envelopes in two hours, and Paula can address 60 envelopes in three hours. About how long will it take them to address 410 envelopes?

OpenStudy (anonymous):

solve for how many envelopes each addresses in an hour Casey's rate = 21/1= 21 Anne's rate = 47/2 = 23.5 Paula's rate = 60/3 = 20 Now if you add there individual rates you get the number of envelopes all three address in one hour. 21 + 23.5 + 20 = 64.5 Now take the number of envelopes you need addressed and divide by the rate the envelopes are addressed at. 410/64.5 = 6.35. . . hours is how long it will take all three to address 410 envelopes (I used . . . because the decimal does not end there)